OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation.

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] that is the formula for point slope

OpenStudy (anonymous):

you do \[8-3/ 6-2= 5/\] so 5/4 is slope

OpenStudy (anonymous):

y-3=5/4(x-2) OR y-8=5/4(x-6)

OpenStudy (anonymous):

there you go did you understand

OpenStudy (anonymous):

Yeah. For the most part. Thanks for all your help!

OpenStudy (anonymous):

no problem